OpenStudy (anonymous):

giving medal for answer... A triangle has an area of 36 m^2 and a base of 6 m. What is the height of the triangle? a. 3 m c. 216 m b. 12 m d. 40 m

OpenStudy (chmvijay):

area of traingle = 1/2 base 8 height 36= 1/2 *6 *height 36*2 /6 = height

OpenStudy (anonymous):

12

OpenStudy (chmvijay):

yup :)